A 12.3 cm piece of copper tube has a mass of 36.50 g. What is the length of tube if it has a mass of 110 kg?











1
Expert's answer
2022-06-14T09:29:04-0400

12,3 cm — 36,5 g

X — 110* 103 g


X = 12,3 * 110 / 36,5 = 37,07*103 cm =

= 370,7 m

Answer: 370,7 m